I have added an October spectrum to the noise budget, it's the first plot below. Click here for the Bokeh version. Sorry it's not on nodus yet.
The Marconi actuator strength was at 10 kHz/Vpp. The PLL OLG used is shown in a plot below. The phase locked loop open loop gain unity gain frequency (PLL OLG UGF) is ~30 kHz.
Not much has changed in the last two months, as we've been running around like chickens with their heads cut off accomplishing little in the way of science but much in the way of making that future science possible. Still, the hard stuff lies in front of us.
Getting down to the predicted coatings thermal noise must be our top priority for the next two months.
This means we need to understand what limits us, i.e. make additions to this noise budget where the total expected and measured agree in some sense. One curve that is obviously wrong is the PLL osc. noise curve in grey. At 10 kHz/Vpp we are ~15 times worse than that curve indicates on this noise budget.
FSS boxes must be better understood to remove suspected low frequency noise injection and avoid saturation of the high gain EOM paths.
EOMs must be looked at to see if thermally stable and the EOM drivers are okay.
Thermal drift of the cavity MUST be understood and eliminated. Perhaps the different laser power resonating in the cavities is causing this. Measurements are currently underway.
Think about adding PMCs or venting, but we must understand WHY we need to do these time-consuming things before we do them. Also good if we can model our expected improvement. If we can get coatings thermal noise without this, all the better.